e. The price of Dwayne Corporation stock is expected to be RM68 in 5 years. Dividends are anticipated to increase at an annual rate of 10 percent from the most recent dividend of RM1.00. If your required rate of return is 15 percent, how much are you willing to pay for Dwayne stock now?
Answer #1

From Dividend discount model with constant growth rate

P = D(1)*/r-g

where D(1) - dividend after one period = D(0)*(1+g)=$1*(1+0,.1) =$1.1

r=15% = 0.15 , g = 10% = 0.1

P = 1.1/(0.15-0.1) = $22